OpenStudy (anonymous):

Line AB is the diameter of a circle with center P. The coordinates of the endpoints of the diameter are (1, 3) and (6, 6). What are the coordinates of P? Possible answers: (2, 6) (2.2, 3) (2.5, 3.5) (4.5, 5) (3.5, 4.5)

OpenStudy (johnweldon1993):

Okay if AB is a diameter that goes ACROSS the circle...that means that at the middle of that diameter...we would have the center of the circle... So we need o find the midpoint of that diameter... We have the 2 endpoints of that to be (1 , 3) (6 , 6) \[\large Midpoint = (\frac{x_1 + x_2}{2},\frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

I'm not sure of what the answer is.

OpenStudy (anonymous):

it is b

OpenStudy (johnweldon1993):

Well since we have the 2 points... (1 , 3) and (6 , 6) We have (x1, y1) and (x2 , y2) So we plug those in \[\large (\frac{1 + 6}{2}, \frac{3 + 6}{2}) = (\frac{7}{2},\frac{9}{2}) = ?\]

OpenStudy (anonymous):

3.5,4.5 I believe

OpenStudy (anonymous):

yea u r right

OpenStudy (johnweldon1993):

Correct @ikileyxx :)

OpenStudy (anonymous):

Thank you I appreciate it!

OpenStudy (johnweldon1993):

Anytime :)